OpenStudy (anonymous):

1. When solving systems of equations, how do you determine what method to use? 2. solve the system and show your steps 2y = x + 1 –2x – y = 7

OpenStudy (abb0t):

You can use elimination or substitution.

OpenStudy (abb0t):

For substitution, you can divide the first equation by 2 get your Y and substitute it.

OpenStudy (abb0t):

\[y = \frac{ x }{ 2 } + \frac{ 1 }{ 2 }\] now you have a y, that you can use to plug in for y for the 2nd equation.

OpenStudy (abb0t):

For elimination, you do the same method, except you are elimiating ONE variable to get just ONE, like for instance, eliminate y and get only x's. but you must have it in the form: 1. (x) + (y) = number 2. (x) - (y) = number notice how you can eliminate Y if you add both together to get: 2x = number then divide by 2 to get x. Then plug it back in and solve for Y! :)

OpenStudy (abb0t):

You can also use gauss elimination and put it into a matrix